# Symmetric Iterations with Countable Support: A Framework for Choiceless ZF Extensions

### v1

@@ 3.6 Iteration length and class-length versions
The successor step is the usual two-step construction. At limits, we use the bookkeeping convention of Remark 2.7: if $\mathrm{cf}(\lambda)>\omega$ we take direct limits, and if $\mathrm{cf}(\lambda)=\omega$ we take the tree (inverse-limit) presentation. In both cases, the limit filter is generated by head-pushforwards together with tuple-stabilizers, and is therefore normal and $\omega_{1}$-complete by Theorem 3.15. The symmetric model at stage $\lambda$ then satisfies $\mathrm{ZF}$ by Theorem 3.20; and if the ambient ground is $\mathrm{ZFC}$, it satisfies $\mathrm{DC}$ by Theorem 3.23. ∎
Theorem 3.27 (Class-length countable-support iterations).
Work in a ground model of $\mathrm{GBC}$ with Global Choice and sufficient class-recursion to carry out the definable iteration (e.g. $\mathrm{ETR}$ for the relevant formulas). Suppose the iteration template is first-order definable over the ground. Then the recursion in Definition 3.1 can be carried out through $\mathrm{Ord}$, yielding a class-length countable-support symmetric iteration and a class forcing $\mathbb{P}_{\mathrm{Ord}}$.
If $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ is $V$-generic and $\mathrm{HS}_{\mathrm{Ord}}$ denotes the class of hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-names with respect to the induced class group and limit filter, then
$M\ :=\ \{\dot{x}^{G}:\dot{x}\in\mathrm{HS}_{\mathrm{Ord}}\}$
is a transitive model of $\mathrm{ZF}$. If the ground satisfies $\mathrm{ZFC}$, then $M\models\mathrm{DC}$.
@@
The definability hypothesis and class-recursion yield the class iteration and the induced class symmetric system. Define $M$ by evaluating hereditarily symmetric class-names, as in the set-length case. The standard symmetric-model verification of the $\mathrm{ZF}$ axioms (cf. Theorem 3.20) relativizes to this class setting: closure of $\mathrm{HS}_{\mathrm{Ord}}$ under the operations needed for Pairing/Union/Separation/Replacement follows from normality and $\omega_{1}$–completeness of the limit filters at the relevant stages, together with the forcing theorem in the $\mathrm{GBC}$ background. The $\mathrm{DC}$ argument likewise relativizes (cf. Theorem 3.23) using countable-tuple closure of $\mathrm{HS}_{\mathrm{Ord}}$. ∎
Remark 3.28 (Formalization and hypotheses).
The $\mathrm{GBC}$ assumption is used only to formalize the class recursion and to speak cleanly about class-length iterations. Global Choice is used in the ambient bookkeeping for definability and class recursion; it is not asserted to hold in $M$.
4 Application Methodology and Comparative Analysis
We now outline the general strategy for employing the framework developed in Section 3.2 and Section 3.4. Throughout, recall that hereditarily symmetric (HS) names are defined via stabilizers of names (Definition 3.16), and that the limit filter $\tilde{\mathcal{F}}_{\lambda}$ is always constructed to be normal and $\omega_{1}$-complete (Theorem 3.15).
We rely on the HS-closure results from Section 3.3 (Theorem 3.17) and the ZF/DC preservation theorems from Sections 3.43.5.
4.1 Employing the framework
To construct a specific model using this framework, one proceeds as follows:
@@ 4.1 Employing the framework
    Specify the Successor System: At each stage $\alpha$, select a symmetric system $(\mathbb{Q}_{\alpha},\mathcal{G}_{\alpha},\mathcal{F}_{\alpha})$.
@@
    Ensure Filter Robustness: A critical requirement of this countable-support framework is that the symmetry filters must remain proper (non-trivial) after being closed under countable intersections. Unlike finite-support iterations, where filters generated by finite stabilizers suffice, this framework requires filters generated by stabilizers of countable sets (or similarly robust generators) to survive the $\omega_{1}$-completion process at limit stages without collapsing the symmetric model to the full generic extension.
@@
    Analyze the Limit Model: Use the preservation theorems (Theorem 3.20 and Theorem 3.23) to guarantee $\mathrm{ZF}+\mathrm{DC}$ in the result. The specific failure of the Axiom of Choice (e.g., the negation of $\mathrm{AC}_{\omega_{1}}$ or the existence of amorphous sets) will depend entirely on the specific combinatorics of the groups $\mathcal{G}_{\alpha}$ and filters $\mathcal{F}_{\alpha}$ chosen in Step 1.
We intentionally defer the construction of specific worked examples to a companion application paper. The landscape of potential applications—ranging from separating specific weak choice principles to investigating partition relations in models of $\mathrm{ZF}+\mathrm{DC}$—is too diverse to treat adequately within the scope of this foundational framework. Furthermore, a detailed analysis of the combinatorial properties of these specific models would significantly expand the length of the present work.
Remark 4.1 (Why finite support does not suffice here).
(1) Countable sequences in the symmetric model. The proof of $\mathrm{DC}$ preservation (Theorem 3.23) does not rely on the forcing partial order being $\omega$-closed. Indeed, this preservation holds even if the iterands are c.c.c. notions (like Cohen forcing) which lack such closure. Instead, the proof relies on the $\omega_{1}$–completeness of the limit filter combined with the closure of $\mathrm{HS}$ names under countable tuples (Theorem 3.17). Any $\omega$-sequence of elements of $M$ arising in $V[G]$ is represented by a countably supported tuple of $\mathrm{HS}$-names, and thus belongs to $M$. Finite-support filters generally lack countable-intersection closure at limits of cofinality $\omega$, and therefore fail to secure the name closure required for $\mathrm{DC}$.
(2) Closure of symmetry under countable tuples. Even if the limit forcing admitted a condition, standard finite-support iterations yield limit filters closed only under finite intersections. Our arguments repeatedly rely on the fact that if a sequence of names $\langle\dot{x}_{n}:n<\omega\rangle$ consists of HS names, then the tuple itself is HS. This requires:
$\bigcap_{n<\omega}\mathrm{fix}(\dot{x}_{n})\in\tilde{\mathcal{F}}_{\lambda}.$
This is exactly the $\omega_{1}$-completeness provided by Theorem 3.15 (and by Remark 3.3 at $\mathrm{cf}(\lambda)\geq\omega_{1}$).
Thus, finite support fails both at the level of conditions (lack of fusion/support for limits) and at the level of symmetry (lack of countable stabilizer intersection). Consequently, the ZF/DC preservation method presented here is not available in the finite-support setting.
@@ 5 Conclusion
We have presented a uniform framework for iterated symmetric extensions with countable support. Our construction handles limit stages of uncountable cofinality via direct limits, while limits of cofinality $\omega$ are handled by a tree (inverse-limit) presentation. A central feature of the construction is that the resulting limit filters are normal and $\omega_{1}$-complete (Theorem 3.15).
This filter completeness is the technical engine of the framework: it yields the closure of hereditarily symmetric names under countable tuples, ensuring the preservation of $\mathrm{ZF}$ (Theorem 3.20). Furthermore, whenever the ambient ground model satisfies $\mathrm{ZFC}$, this same completeness ensures that the resulting symmetric model satisfies $\mathrm{DC}$ (Theorem 3.23).
This toolkit is designed to be reusable. It provides the machinery for both set-length and class-length iterations (Theorem 3.27), allowing future application papers to focus on the specific combinatorial design of iterands and symmetry filters needed to obtain specific failures of choice, while citing the present work for the underlying ZF and DC preservation machinery.
@@
AI tools. Portions of the exposition were drafted and revised with assistance from large language models (LLMs), which were used as a writing aid for phrasing, organization, typesetting, and error-checking suggestions. All mathematical statements, proofs, and final editorial decisions were made by the author.

### v3

@@ 3.6 Iteration length
The successor step is the standard two-step symmetric-system construction (as cited in Definition 3.1). At limit stages $\lambda\leq\Theta$, we use Remark 3.2 to identify the countable-support limit with the direct limit when $\mathrm{cf}(\lambda)\geq\omega_{1}$, and we use the usual countable-support definition when $\mathrm{cf}(\lambda)=\omega$. In both cofinality cases, the limit filter is defined by Definition 3.9 and is normal and $\omega_{1}$-complete by Theorem 3.13. The symmetric model at stage $\lambda$ satisfies $\mathrm{ZF}$ by Theorem 3.17; and if the ambient ground is $\mathrm{ZFC}$, it satisfies $\mathrm{DC}$ by Theorem 3.20. ∎
Remark 3.27 (Class-length iterations).
We do not claim a general class-length $\mathrm{ZF}$-preservation theorem in this paper. The set-length results (Proposition 3.26, Theorems 3.17 and 3.20) are stated and proved for set-length iterations only, where the ambient forcing extension satisfies $\mathrm{ZFC}$ automatically.
At class length, $\mathrm{ZF}$ preservation for symmetric submodels requires hypotheses beyond pretameness: there are class forcings that are pretame yet whose symmetric submodels fail Collection or Power Set. Identifying the correct general hypothesis is a delicate open problem in class forcing theory.
Specific class-length symmetric iterations can often be analyzed directly by verifying each $\mathrm{ZF}$ axiom within the particular construction, as in [5]. The present framework supplies the limit-filter technology ($\omega_{1}$-completeness, Theorem 3.13) that such a verification would cite when checking closure of hereditarily symmetric names under countable tuples. We leave the formulation of a general class-length theorem to future work.
4 Application: The Iterated Unordered Pairs Model
We illustrate the framework with a self-contained construction that, for any uncountable cardinal $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$, produces a model of $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$, where $\mathcal{F}$ is a $\kappa$-indexed family of $2$-element sets of reals carrying no choice function. Each step of the iteration adds exactly one new unpickable pair, so the degree of $\mathrm{AC}$-failure is dialed by the iteration length. We also prove that the natural finite-support analogue fails to preserve $\mathrm{DC}$ at the first limit stage, showing that countable support is not merely sufficient but structurally necessary.
4.1 Setup: the step forcing and its symmetry
Fix an uncountable cardinal $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$ and a ground model $V\models\mathrm{ZFC}$. For each ordinal $\alpha<\kappa$ the step forcing is
$\dot{\mathbb{Q}}_{\alpha}\;=\;\mathrm{Add}(\omega)\times\mathrm{Add}(\omega),$
adding two independent Cohen reals, with the group $\dot{\mathcal{H}}_{\alpha}=\mathbb{Z}/2\mathbb{Z}$ acting by swapping coordinates: the nontrivial element $\sigma_{\alpha}$ sends $(s,t)\mapsto(t,s)$. Let $\dot{\mathcal{K}}_{\alpha}$ be the principal filter on $\dot{\mathcal{H}}_{\alpha}=\mathbb{Z}/2\mathbb{Z}$ generated by the whole group: the only filter element is $\mathbb{Z}/2\mathbb{Z}$ itself, i.e. $K\leq\mathbb{Z}/2\mathbb{Z}$ belongs to $\dot{\mathcal{K}}_{\alpha}$ if and only if $K=\mathbb{Z}/2\mathbb{Z}$.
After the generic at stage $\alpha$ is added, let $a_{\alpha},b_{\alpha}$ denote the two resulting Cohen reals. The individual reals $a_{\alpha}$ and $b_{\alpha}$ are not HS names: $\operatorname{sym}(\dot{a}_{\alpha})=\{\mathrm{id}\}\notin\dot{\mathcal{K}}_{\alpha}$, since $\{\mathrm{id}\}\neq\mathbb{Z}/2\mathbb{Z}$. The unordered pair
$\dot{P}_{\alpha}\;:=\;\{\dot{a}_{\alpha},\,\dot{b}_{\alpha}\}$
is an HS name: $\operatorname{sym}(\dot{P}_{\alpha})=\mathbb{Z}/2\mathbb{Z}\in\dot{\mathcal{K}}_{\alpha}$, because $\sigma_{\alpha}$ merely swaps the two elements of the pair and so fixes the pair as a set. Set
$\dot{\mathcal{F}}\;:=\;\bigl\langle\dot{P}_{\alpha}:\alpha<\kappa\bigr\rangle,$
the $\mathbb{P}_{\kappa}$-name for the function $\alpha\mapsto P_{\alpha}$, a $\kappa$-indexed family of $2$-element sets of reals.
Lemma 4.1 (Automorphisms used in the unordered-pairs argument).
In the countable-support iteration of $\kappa$ steps described above:
  1. 1.
    $\mathcal{G}_{\kappa}$ is abelian.
  2. 2.
    For each $\alpha<\kappa$, there exists an automorphism $g_{\alpha}\in\mathcal{G}_{\kappa}$ which acts by the nontrivial swap $\sigma_{\alpha}$ at stage $\alpha$ and acts trivially at every stage $\gamma\neq\alpha$.
  3. 3.
    For every $\beta\leq\alpha<\kappa$,
    $\rho_{\beta,\kappa}(g_{\alpha})=\mathrm{id}.$
    In particular, $g_{\alpha}\in\ker(\rho_{\beta,\kappa})$ whenever $\beta\leq\alpha$.
@@ 4.1 Setup: the step forcing and its symmetry
For (1), argue by induction on $\delta\leq\kappa$ that $\mathcal{G}_{\delta}$ is abelian. At stage 0 this is trivial. At a successor stage, the new step group is $\mathbb{Z}/2\mathbb{Z}$, which is abelian, and the successor-stage action is coordinatewise, so automorphisms commute stage by stage. At a countable-cofinality limit, $\mathcal{G}_{\lambda}$ is an inverse limit of the earlier $\mathcal{G}_{\beta}$, hence abelian. At an uncountable-cofinality limit, $\mathcal{G}_{\lambda}$ is identified with $\bigcup_{\beta<\lambda}\mathcal{G}_{\beta}$ by Definition 3.1, hence again abelian.
For (2), fix $\alpha<\kappa$. At stage $\alpha+1$, the step group contributes the nontrivial swap $\sigma_{\alpha}$ exchanging the two Cohen coordinates added at stage $\alpha$. Extending by the identity at every later stage yields a coherent family of automorphisms, hence an element $g_{\alpha}\in\mathcal{G}_{\kappa}$ acting nontrivially only at stage $\alpha$.
For (3), if $\beta\leq\alpha$, then by construction $g_{\alpha}$ is trivial on every stage below $\beta$, so its restriction to stage $\beta$ is the identity. This is exactly the statement $\rho_{\beta,\kappa}(g_{\alpha})=\mathrm{id}$. ∎
4.2 Main theorem
Theorem 4.2 (Iterated Unordered Pairs Model).
Let $\kappa$ be an uncountable cardinal with $\mathrm{cf}(\kappa)\geq\omega_{1}$, and let $M_{\kappa}$ be the symmetric model obtained by the $\kappa$-step countable-support symmetric iteration described above. Then:
@@ 4.2 Main theorem
    $M_{\kappa}\models\mathrm{ZF}+\mathrm{DC}$.
@@
    The indexed family $\mathcal{F}=\langle P_{\alpha}:\alpha<\kappa\rangle\in M_{\kappa}$ has no choice function in $M_{\kappa}$.
@@
    $M_{\kappa}\models\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$.
Proof.
(1) $\mathrm{ZF}+\mathrm{DC}$. At each limit stage $\lambda\leq\kappa$, Theorem 3.13 gives a normal $\omega_{1}$-complete limit filter $\tilde{\mathcal{F}}_{\lambda}$. $\mathrm{ZF}$ preservation is Theorem 3.17. Since $\kappa$ is a set ordinal and $V\models\mathrm{ZFC}$, $\mathrm{DC}$ preservation is Theorem 3.20.
[…]
$\operatorname{sym}(\dot{\mathcal{F}})=\mathcal{G}_{\kappa}\in\tilde{\mathcal{F}}_{\kappa}.$
[…]
Now suppose for contradiction that $f\in M_{\kappa}$ is a choice function for $\mathcal{F}$, i.e. $f(\alpha)\in P_{\alpha}=\{a_{\alpha},b_{\alpha}\}$ for each $\alpha<\kappa$. Let $\dot{f}$ be an HS name for $f$, so $\operatorname{sym}(\dot{f})\in\tilde{\mathcal{F}}_{\kappa}$.
Since $\mathcal{G}_{\kappa}$ is abelian (Lemma 4.1(1)), conjugation in Definition 3.9(1) is vacuous. Hence membership $\operatorname{sym}(\dot{f})\in\tilde{\mathcal{F}}_{\kappa}$ is witnessed by finitely many head pullbacks: there exist $m<\omega$, stages $\beta_{0},\dots,\beta_{m-1}<\kappa$, and subgroups $H_{i}\in\mathcal{F}_{\beta_{i}}$ such that
$\bigcap_{i<m}\rho_{\beta_{i},\kappa}^{-1}(H_{i})\ \leq\ \operatorname{sym}(\dot{f}).$
Let $\beta^{*}=\max_{i<m}\beta_{i}<\kappa$. For any $\alpha\geq\beta^{*}$, Lemma 4.1(3) gives $g_{\alpha}\in\ker(\rho_{\beta_{i},\kappa})\leq\rho_{\beta_{i},\kappa}^{-1}(H_{i})$ for each $i<m$, hence $g_{\alpha}\in\operatorname{sym}(\dot{f})$.
Since $g_{\alpha}\in\operatorname{sym}(\dot{f})$, we have $g_{\alpha}(\dot{f})=\dot{f}$, so in the symmetric extension $g_{\alpha}(f(\alpha))=f(\alpha)$. But by Lemma 4.1(2), $g_{\alpha}$ applies the nontrivial swap at stage $\alpha$, exchanging $a_{\alpha}$ and $b_{\alpha}$: if $f(\alpha)=a_{\alpha}$, then $g_{\alpha}(a_{\alpha})=b_{\alpha}\neq a_{\alpha}$, and similarly if $f(\alpha)=b_{\alpha}$. This contradiction shows that no choice function for $\mathcal{F}$ exists in $M_{\kappa}$.
(3) Combines (1) and (2). ∎
Remark 4.3 (Each step does something; the length is the dial).
The symmetry argument in part (2) applies to each $\alpha$ independently: the $\mathbb{Z}/2\mathbb{Z}$ swap at stage $\alpha$ makes the pair $P_{\alpha}$ unpickable regardless of what happens at other stages. Running the iteration to any uncountable cardinal $\kappa^{\prime}\leq\kappa$ with $\mathrm{cf}(\kappa^{\prime})\geq\omega_{1}$ produces a model in which the truncated indexed family
$\langle P_{\alpha}:\alpha<\kappa^{\prime}\rangle$
has no choice function. By choosing $\kappa$ freely — any uncountable cardinal with $\mathrm{cf}(\kappa)\geq\omega_{1}$, with $\omega_{1}$ itself the simplest instance — one obtains $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$ at any such level. The restriction $\mathrm{cf}(\kappa)\geq\omega_{1}$ is used in the proof of part (2): the filter-membership witness for $\operatorname{sym}(\dot{f})$ then has finitely many head-pullback stages, all bounded below $\kappa$ by a single $\beta^{*}<\kappa$, allowing the tail-flip argument to go through. The case $\mathrm{cf}(\kappa)=\omega$ requires a separate argument tracking the cofinal sequence of witness stages and is not treated here.
4.3 Why finite support fails here
Proposition 4.4 (Finite support cannot preserve $\mathrm{DC}$ at the $\omega$-th limit).
Suppose one applies the same step forcing at each stage $\alpha<\omega$—that is, $\dot{\mathbb{Q}}_{\alpha}=\mathrm{Add}(\omega)\times\mathrm{Add}(\omega)$ with $\mathbb{Z}/2\mathbb{Z}$ symmetry—but uses a finite-support symmetric iteration. Then the resulting symmetric model $M^{\mathrm{fs}}_{\omega}$ fails $\mathrm{DC}$.
Proof.
Let $A:=\bigcup_{n<\omega}P_{n}$ be the union of all pairs in $M^{\mathrm{fs}}_{\omega}$ and define a binary relation on $A$ by
$x\mathrel{R}y\quad\Longleftrightarrow\quad\exists n<\omega\;(x\in P_{n}\;\wedge\;y\in P_{n+1}).$
The relation $R$ is total ($\forall x\in A\;\exists y\in A\;x\,R\,y$) and belongs to $M^{\mathrm{fs}}_{\omega}$ (its definition uses only the HS names $P_{n}$ and the successor function on $\omega$, both symmetric).
A $\mathrm{DC}$-sequence for $(A,R)$ would be an $\omega$-sequence $\langle x_{n}:n<\omega\rangle$ with $x_{n}\in P_{n}$ for each $n$, i.e., a partial choice function for $\{P_{n}:n<\omega\}$. Suppose $\dot{f}$ is an HS name for such a sequence in $M^{\mathrm{fs}}_{\omega}$. Then $\operatorname{sym}(\dot{f})$ lies in the finite-support limit filter $\tilde{\mathcal{F}}^{\mathrm{fs}}_{\omega}$ at stage $\omega$.
In the finite-support setting, every filter element at stage $\omega$ contains a subgroup that fixes all but finitely many coordinates: it is generated by finitely many head pullbacks $\rho_{k,\omega}^{-1}(H_{k})$ for $k<m$ and $H_{k}\in\mathcal{F}_{k}$. Hence for all but finitely many $n<\omega$, the swap automorphism $g_{n}$ (applying $\sigma_{n}$ at stage $n$, trivial elsewhere) belongs to $\operatorname{sym}(\dot{f})$.
But if $g_{n}\in\operatorname{sym}(\dot{f})$ then $g_{n}$ fixes $\dot{f}$, so $g_{n}(f(n))=f(n)$. Since $g_{n}$ swaps the two elements of $P_{n}$, this forces $f(n)$ to be fixed by the swap $a_{n}\leftrightarrow b_{n}$—impossible, as neither $a_{n}$ nor $b_{n}$ is individually HS. Contradiction; so no HS name for a DC-sequence through $(A,R)$ exists, and $M^{\mathrm{fs}}_{\omega}\not\models\mathrm{DC}$. ∎
Remark 4.5 (Why countable support resolves this).
In the countable-support iteration the limit filter $\tilde{\mathcal{F}}_{\omega}$ at stage $\omega$ is the smallest normal $\omega_{1}$-complete filter extending the head-generator family $\mathcal{B}_{\omega}$ (Definition 3.9(2) and Theorem 3.13). An HS name $\dot{f}$ for the DC-sequence only needs $\operatorname{sym}(\dot{f})\supseteq\bigcap_{n<\omega}\operatorname{sym}(\dot{x}_{n})$, and $\omega_{1}$-completeness guarantees this intersection lies in $\tilde{\mathcal{F}}_{\omega}$ (Theorem 3.15). No individual swap $\sigma_{n}$ need belong to $\operatorname{sym}(\dot{f})$; the tuple name is hereditarily symmetric as a whole, while each $P_{n}$ remains unpickable for the same reason as before. This is the precise sense in which countable support is not a convenience but a structural requirement.
@@ 5 Conclusion
We isolated the limit-stage filter construction needed for countable-support symmetric iterations built from the standard successor-step theory. The resulting limit filters are canonical, normal, and $\omega_{1}$-complete (Theorem 3.13). This $\omega_{1}$-completeness is the key input for closure of hereditarily symmetric names under countable tuples (Theorem 3.15), for the standard verification of $\mathrm{ZF}$ (Theorem 3.17), and, over a $\mathrm{ZFC}$ ground, for $\mathrm{DC}=\mathrm{DC}_{\omega}$ in the resulting symmetric model (Theorem 3.20).
We do not claim a general class-length $\mathrm{ZF}$-preservation theorem here. Specific class-length symmetric iterations may still be analyzable separately, but such results require additional hypotheses and care beyond the present set-length framework.
As a self-contained illustration, the iterated unordered pairs model (Section 4) demonstrates that for any uncountable cardinal $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$ the framework produces a model of $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$, with the degree of $\mathrm{AC}$-failure dialed by the iteration length. The same construction proves that finite support is structurally insufficient: the natural finite-support analogue fails $\mathrm{DC}$ at the first $\omega$-limit stage (Proposition 4.4).
In this sense the paper can be used modularly: once a concrete construction specifies its successor-stage symmetric systems, the limit-filter and $\mathrm{ZF}/\mathrm{DC}$ preservation arguments developed here may be cited separately, with any additional properties (e.g. chain conditions or cardinal behavior) verified in the specific application.
@@
AI tools. Large language models (LLMs) were used as a writing aid for copy-editing, typesetting, and error-checking suggestions. All mathematical statements, proofs, and final editorial decisions were made by the author with whom final responsibility resides.